OpenStudy (anonymous):

write always, sometimes, or never. If it is always true, give the property or two examples. If it is never true, give the property or two examples. If it is sometimes true, give one example for which it is true and one example for which it is false. x + x = 2x

OpenStudy (anonymous):

Always true, when you add x + x, you can combine the two and get 2x. So it makes the equation 2x = 2x true. You can combine because adding x two times is the same thing as multiplying x by 2. Ex: x = 5 5 + 5 = 2(5) 10 = 10 Examples: 2x + x = 3x 3x + 2x = 5x